OpenStudy (anonymous):

If f(x) = x^2 − 2x + 9 and g(x) = 8 − x, what is (f o g)(−4)?

OpenStudy (anonymous):

Select one of the options below as your answer: A. 111 B. 144 C. 120 D. 129

OpenStudy (anonymous):

\[g(-4)=8-(-4)=12\] \[f(12)=12^2-2\times 12+9=129\] \[